# Computing the Elementary Symmetric Polynomials in Positive Characteristics

### v2

@@ Abstract
We first extend the results of Chatterjee, Kumar, Shi, Volk (Computational Complexity 2022) by showing that the degree $d$ elementary symmetric polynomials in $n$ variables have formula lower bounds of $\Omega(d(n-d))$ over fields of positive characteristic. Then, we show that the results of the universality of the symmetric model from Shpilka (Journal of Computer and System Sciences 2002) and the results about border fan-in two $\Sigma\Pi\Sigma$ circuits from Kumar (ACM Trans. Comput. Theory 2020) over zero characteristic fields do not extend to fields of positive characteristic. In particular, we show that
@@
    There are polynomials that cannot be represented as linear projections of the elementary symmetric polynomials (in fact, we show that they cannot be represented as the sum of $k$ such projections for a fixed $k$) and
@@
To prove the first result, we consider a geometric property of the elementary symmetric polynomials, namely, the set of all points in which the polynomial and all of its first-order partial derivatives vanish. It was shown in Meckler, Zaimi and Limaye, Mittal, Pareek that the dimension of this space was exactly d2d-2 for fields of zero characteristic. We extend this to fields of positive characteristic by showing that this dimension must be between d2d-2 and d1d-1. In fact, we show this bound is tight, in the sense that there are (infinitely many) polynomials where each of these bounds is exact.
Then, to consider the border top fan-in of the symmetric model and depth-$3$ circuits (sometimes called border affine Chow rank), we show that it is sufficient to consider the border top fan-in of the sum of linear projections of the elementary symmetric polynomials. This is done by constructing an explicit metapolynomial to check the condition, meaning that this result also applies in the border setting.
@@ 1 Introduction
Given $n$ independent variables $x_{1},\ldots,x_{n}$ and a degree $d\leq n$, we define the elementary symmetric polynomial of degree $d$ (over an arbitrary field $\mathbb{F}$ of characteristic denoted by $\operatorname{char}(\mathbb{F})$) to be the sum of all degree $d$ multilinear monomials in these indeterminants, i.e.,
$e_{n}^{d}=\sum_{S\in{[n]\choose d}}\prod_{i\in S}x_{i}.$
@@
This polynomial naturally appears in many problems, for instance, if we consider the product of affine, constant-free polynomials, we can express such a product by
$\prod_{i=1}^{n}(x_{i}+y)=\sum_{k=0}^{n}y^{n-k}e_{k}^{n}(x_{1},\ldots,x_{n}).$ (1)
@@
We say that these polynomials are symmetric because they are invariant under permutations of the variables, and we say that they are elementary because they are the basic building blocks of symmetric polynomials, namely, the set of symmetric polynomials can be written by $\mathbb{F}[e_{0}^{n},\ldots,e_{n}^{n}]$. We include more fundamental properties in the appendix, Section A.
Motivation.
@@
In algebraic complexity, computation of the elementary symmetric polynomials appear very naturally in many contexts. Algebraically, we commonly define computation through algebraic circuits, which we model using acyclic directed graphs, where certain nodes represent operations, addition and multiplication, and other nodes represent inputs, which are given as independent variables and field constants. We can then restrict this to so-called algebraic formulas, which require that the underlying graph be a tree. We can then further restrict the depth of a tree and force layers to alternate in operations, creating, for example, depth-three $\Sigma\Pi\Sigma$ formulas, where the root node is an addition node. We may further restrict this model to $\Sigma^{[k]}\Pi\Sigma$ formulas, where the root node may only have at most $k$ children.
The earliest-known motivation for studying the complexity of the elementary symmetric polynomials came from boolean circuit complexity, where circuits compute boolean functions using a graph labeled with and-gates, or-gates, and not-gates. When considering boolean formulas of constant depth, it was shown in [FSS84] that the majority function had super-polynomial lower bounds, where the majority function is given by $\bigvee_{I\in{[n]\choose n/2}}\bigwedge_{i\in I}x_{i}$. Clearly, $e_{n,n/2}$ looks like an algebraic analogue of the majority function, so it was believed to be a good candidate for a polynomial that would have constant-depth formulas with super-polynomial lower bounds. Unfortunately, due to a construction of Ben-Or (see Theorem 3.1 of [Shp02]), $e_{n,d}$ was shown to have $\Sigma\Pi\Sigma$ formulas of size $O(n^{2})$ (but the problem of super-polynomial $\Sigma\Pi\Sigma$ lower bounds has since been solved, see [LST24]). In [SW01] and [Shp02], it was shown that this upper bound is tight for $\Sigma\Pi\Sigma$ formulas of certain elementary symmetric polynomials over fields of characteristic zero. This was further extended in [CKSV22] to be tight over general algebraic formulas, also over fields of characteristic zero. We further have that [HY11] found super-polynomial lower bounds for certain elementary symmetric polynomials for homogeneous multilinear formulas, where each node computes a homogeneous multilinear polynomial (the result of which is independent of field characteristic).
@@
Another motivation was explored in [LST24] and [FLST24], where we can view the computation of the elementary symmetric polynomials as naturally related to constant-depth arithmetic formulas. For example, we can consider $\Sigma\Pi\Sigma$ formulas and explore how they can be converted into homogeneous $\Sigma\Pi\Sigma$ formulas, where each node computes a homogeneous polynomial. One idea is to apply (1) to each product gate and then compute each resulting elementary symmetric polynomial using homogeneous circuits (for instance, we could use the upper bound of [SW01] for homogeneous computation of $e_{d}^{n}$ by $\Pi\Sigma\Pi\Sigma$ of size $n2^{O(\sqrt{d})}$, but this only works for characteristic zero fields). This idea can be extended to work for deeper constant-depth circuits.
In [Shp02], the elementary symmetric polynomials were used to define an algebraic computational model, which was called the symmetric model. It was defined by taking a series of linear polynomials $L_{1},\ldots,L_{m}\in\mathbb{F}[x_{1},\ldots,x_{n}]$ and a degree $d\leq m$ and defining the computation by $e_{d}^{m}(L_{1},\ldots,L_{m})$. [Shp02] then showed that the symmetric model is universal over fields of characteristic zero, meaning that it can compute any (homogeneous) polynomial. This was done using Fischer’s identity (see [Fis94] and [Shp02]), which tells us that we can write any degree $d$ homogeneous polynomial $f\in\mathbb{F}[x_{1},\ldots,x_{n}]$ as $f=L_{1}^{d}+\dots+L_{m}^{d}$, for some linear polynomials $L_{1},\ldots,L_{m}\in\mathbb{F}[x_{1},\ldots,x_{n}]$, and it was then shown that polynomials written in this form can be expressed in the symmetric model. We observe that this only works over fields of characteristic zero, as there are many polynomials in positive characteristic that cannot be written in this way, as, for example, $(x+y)^{p}=x^{p}+y^{p}$ when $\operatorname{char}(\mathbb{F})=p$.
Then, in [Kum20], the results around the symmetric model were used to study the border “affine Chow rank” of a polynomial, where this name selected based on [Dut25]. We say that the affine Chow rank of a polynomial $f\in\mathbb{F}[x_{1},\ldots,x_{n}]$ is the smallest $k$ such that $f$ can be computed by a $\Sigma^{[k]}\Pi\Sigma$ circuit. We mention that there are polynomials with affine Chow rank that is $\Omega(n)$ over any field. Recall that, given a complexity measure, we define its border complexity by extending the underlying field to $\mathbb{F}(\epsilon)$, for some new variable $\epsilon$, and, instead of requiring our model to compute $f(x)$, we need only to compute $\epsilon^{N}\cdot f+\epsilon^{N+1}\cdot F(x,\epsilon)$ for some $N\geq 0$ and $F\in\mathbb{F}[x_{1},\ldots,x_{n},\epsilon]$ (where we may sometimes write that $\epsilon^{N}\cdot f+\epsilon^{N+1}\cdot F(x,\epsilon)\simeq f$). Through the results of the symmetric model, [Kum20] showed that, over fields of characteristic zero, every homogeneous polynomial has border affine Chow rank of at most two.
In this paper, we will consider these results over fields of positive characterstic. This is an interesting consideration, as, for algebraic complexity theory, the underlying field of our model “does not matter,” in the sense that collapsing VP and VNP over any single field is enough to collapse the polynomial hierarchy. There have been many recent results that attempt to take known results over fields of certain characteristics and extend them to arbitrary fields (see [And20], [DIK+24], [For24], [BLRS25], [BKR+25]).
Results.
@@
We have considered computation of the elementary symmetric polynomials using formulas over fields of positive characteristic. By the Ben-Or construction, we know that there is a $\Sigma\Pi\Sigma$ circuit computing $e_{d}^{n}$ of size $O(n^{2})$ over infinite fields (independent of characteristic). In this paper, we extend the results of [CKSV22] and show that their results on fields of zero characteristic extend to fields of positive characteristic, namely that this upper bound is tight for formulas for certain elementary symmetric polynomials.
@@
In the second part of the paper, we show that the results of [Shp02] and [Kum20] do not extend to fields of positive characteristic. First, we show that the symmetric model is not universal over such fields. In fact, we show something stronger; we show that there are some polynomials that cannot be represented by $k$-linear combinations of projections of the symmetric model, even if we extend to the border setting.
@@
If $\operatorname{char}(\mathbb{F})\neq 0$, then, for every fixed $n$, there exists a homogeneous polynomial $f\in\mathbb{F}[x_{1},\ldots,x_{n}]$ of degree $d$ such that, for every linear $L_{1}^{(1)},\ldots,L_{m_{1}}^{(1)},L_{1}^{(2)},\ldots,L_{m_{k}}^{(k)}\in\mathbb{F}[x_{1},\ldots,x_{n}]$ and constants $c_{1},\ldots,c_{k}\in\mathbb{F}$,
$f\neq\sum_{i=1}^{k}c_{i}e_{d}(L_{1}^{(i)},\ldots,L_{m}^{(i)}).$ (2)
Further, this applies in the border setting.
As we have previously mentioned, the symmetric model is intimately related to the models studied in [Kum20], namely the $\Sigma^{[k]}\Pi\Sigma$ model. Through this, as an immediate corollary of Theorem 1.2, we show that the results of [Kum20] also do not extend to fields of positive characteristic. In fact, we prove, under such bounds, $\Omega(n)$ lower bounds for the border affine Chow rank of certain polynomials.
@@
If $\operatorname{char}(\mathbb{F})\neq 0$, then, for a fixed $n$, there exists a polynomial in at most $n$ variables that is not in $\overline{\Sigma^{[k]}\Pi\Sigma}$ for $k=o(n)$.
Proof techniques.
We will first focus on proving Theorem 1.1 in Section 2. This proof will largely rely on the proof for the case of characteristic zero from [CKSV22] but with a modification to the step that relies on the field characteristic. Namely, the proof revolves around studying what we will call the set of order-two zeros of a polynomial.
Algebraic complexity theory focuses on using the algebraic properties of polynomials to split them into classes based on how “hard” they are to compute. One such property that has been utilized (see [Gat87], [Kum19], [CKSV22], [KV22], [ABV17]) is what we will call the order-$2$ zero space, which represents the zeros of a polynomial of order at least two. Given a polynomial $f\in\mathbb{F}[x_{1},\ldots,x_{n}]$, we define its order-$2$ zero space, denoted $V_{2}(f)$, to be the points where it vanishes along with its first-order partial derivatives, i.e.,
@@
This idea of the order-$2$ zero space is utilized in [Kum19] and [CKSV22] without a name. Then, in [KV22], it is introduced as the “singular locus” of a polynomial, denoted $\operatorname{sing}(f)$, which is a well-known object in algebraic geometry, in a nod to the notation of [Gat87]. For this paper, we have decided to change this notation, as this definition of singular locus does not precisely align with the algebraic geometric definition. Although it is true that, when $f$ is square-free, this definition exactly characterizes the singular locus ($V_{2}(f)=\operatorname{sing}(f)$), this is not true in general. This follows from the fact that the singular locus of a polynomial is a property of its corresponding hypersurface, a purely geometric object, while the order-$2$ zero space is a property of the polynomial, itself. Specifically, if we consider the power polynomial $p_{d}^{n}=x_{1}^{d}+\dots+x_{n}^{d}$, it is well-known that $\operatorname{sing}(p_{d}^{n})=\{0\}$ (see Example 10.21 of [Gat21]), but, if we consider $\operatorname{char}(\mathbb{F})=q\neq 0$, we trivially observe that $V_{2}(p_{q}^{n})=V(p_{q}^{n})$. We find that letting this definition differ from its natural geometric definition is confusing, so we have, therefore, decided to change it.
@@
Lemma 1.5 (See Lemma 1.7 of [Kum19] and Lemma 3.4 of [CKSV22]).
@@
where the first inequality comes from Lemma 5.8 of [CKSV22] and the last inequality is a basic fact from algebraic geometry. ∎
@@
The key observation for the proof in [CKSV22] was that small formulas computed polynomials with “many” order-$2$ zeros. Then, they prove an upper bound of the dimension of the set of order-$2$ zeros of the elementary symmetric polynomials and utilize this to show an $\Omega(d(n-d))$ lower bounds on formulas that compute them over fields of zero characteristic (with $n$ being the number of variables and $d$ being the degree). For certain selections of $n$ and $d$, this bound is tight, as the Ben-Or construction shows an $O(n^{2})$ upper bound on the size of $\Sigma\Pi\Sigma$ formulas computing $e_{d}^{n}$. This was known to be tight for $\Sigma\Pi\Sigma$ formulas over fields of characteristic zero from [Shp02].
It was shown in [MZ17], [LMP19], and [CKSV22] that $\dim V_{2}(e_{d}^{n})=d-2$ over fields of characteristic zero, where dimension is, of course, defined in terms of affine varieties. No such equivalent statement was known (to the author’s knowledge) for fields of positive characteristic. In this paper, we show that the dimension is slightly different over such fields, specifically, it can vary between being d2d-2 or d1d-1.
@@
For $\operatorname{char}(\mathbb{F})=p\neq 0$, we have that $d-2\leq\dim V_{2}(e_{d}^{n})\leq d-1$. In particular, for fixed $d\geq 1$, there are values of $n$ such that $\dim V_{2}(e_{d}^{n})=d-2$ and values such that $\dim V_{2}(e_{d}^{n})=d-1$.
Then, for the rest of the proof in [CKSV22] to work, we need only that $\dim V_{2}(e_{d}^{n})\leq d$, so this lemma suffices for proving the main result.
Interestingly, Lemma 1.7 allows us to also extend the lower bound result in [Shp02] of $e_{d}^{n}$ in $\Sigma\Pi\Sigma$ formulas of $\Omega(d(n-d))$ to fields of positive characteristics. In [Shp02], they show that, over fields of characteristic zero, for every vector space $V\subseteq\mathbb{F}^{n}$ such that $e_{d}^{n}$ vanishes on $V$, we have that $\dim(V)<\frac{n+d}{2}$. For fields of positive characteristic, observe that if we combine Proposition 6 from [GGIL22] with Lemma 1.7, we conclude that we have the upper bound of $\dim(V)\leq\frac{n+d-1}{2}$ [^2] Using the fact that we can represent a polynomial as a “strength” decomposition of size $\operatorname{codim}(V)$.
@@
Then, we clearly have a product decomposition, so we can use Lemma 1.5 to conclude that $\dim V_{2}(f)\geq n-\ell\cdot 2^{d+1}$. Then, if we select $d$ to not depend on $n$, we conclude that $\ell\geq\Omega(n-\dim V_{2}(f))$. We should observe that this argument only works in the non-border case, and it would require a careful argument to show that it works in the border setting.

### v3

@@ Abstract
We first extend the results of Chatterjee, Kumar, Shi, Volk (Computational Complexity 2022) by showing that the degree $d$ elementary symmetric polynomials in $n$ variables have formula lower bounds of $\Omega(d(n-d))$ over fields of positive characteristic. Then, we show that the results of the universality of linear projections of elementary symmetric polynomials from Shpilka (Journal of Computer and System Sciences 2002) and of border fan-in two $\Sigma\Pi\Sigma$ circuits from Kumar (ACM Trans. Comput. Theory 2020) over zero characteristic fields do not extend to fields of positive characteristic. In particular, we show that
@@
    There are polynomials that cannot be represented as linear projections of the elementary symmetric polynomials (in fact, we show linear lower bounds over the size of the sum of such linear projections) and
@@
To prove the first result, we consider a geometric property of the elementary symmetric polynomials, namely, the set of all points in which the polynomial and all of its first-order partial derivatives vanish. It was previously shown that the dimension of this space was exactly d2d-2 for fields of zero characteristic. We extend this to fields of positive characteristic by showing that this dimension must be between d2d-2 and d1d-1. In fact, we provide some criterion where it is d2d-2 and others where it is d1d-1.
Then, to consider the border top fan-in of linear projections of the elementary symmetric polynomials and border depth-$3$ circuits (sometimes called border affine Chow rank), we show that it is sufficient to consider the border top fan-in of the sum of such linear projections of the elementary symmetric polynomials. This is done by an explicit construction of a ’metapolynomial,’ meaning that this result also applies in the border setting.
@@ 1 Introduction
Given $n$ independent variables $x_{1},\ldots,x_{n}$ and a degree $d\leq n$, we define the elementary symmetric polynomial of degree $d$ (over an arbitrary field $\mathbb{F}$ of characteristic denoted by $\operatorname{char}(\mathbb{F})$) to be the sum of all degree $d$ multilinear monomials, i.e.,
$e_{d}^{n}=\sum_{S\in{[n]\choose d}}\prod_{i\in S}x_{i}.$
@@
This polynomial naturally appears in many situations; for instance, they show up when considering the product of affine polynomials.
$\prod_{i=1}^{n}(x+a_{i})=\sum_{k=0}^{n}x^{n-k}e_{k}^{n}(a_{1},\ldots,a_{n}).$ (1)
@@
We say that these polynomials are symmetric because they are invariant under permutations of the variables, and we say that they are elementary because they are the basic building blocks of symmetric polynomials; namely, the set of symmetric polynomials can be expressed as $\mathbb{F}[e_{0}^{n},\ldots,e_{n}^{n}]$. We include more fundamental properties in the appendix, Section A.
Motivation
@@
In this paper, we will study the computation of elementary symmetric polynomials through the lens of algebraic complexity theory. We commonly define computation through algebraic circuits, which we model using acyclic directed graphs, where certain nodes represent the operations of addition or multiplication and other nodes represent inputs, which are given as independent variables and field constants. We can then restrict this to so-called algebraic formulas, which require that the underlying graph be a tree (each node has at most one outgoing edge). We can further restrict the depth of a tree and force layers to alternate in operations, creating, for example, depth-three $\Sigma\Pi\Sigma$ formulas, where we force the root node to be an addition node. We may further restrict this model to $\Sigma^{[k]}\Pi\Sigma$ formulas, where the root node may only have at most $k$ children.
Elementary symmetric polynomials
@@
Elementary symmetric polynomials are very common polynomials that appear naturally in various algebraic complexity problems. These polynomials are relatively easy to compute (they can be computed by $\Sigma\Pi\Sigma$ formulas of size $O(n^{2})$), so there is hope they are simple enough that a deep understanding of them is achievable. One recent motivation of them was explored in [24] and [12], where the computation of the elementary symmetric polynomials naturally appears from studying constant-depth arithmetic formulas. For example, when considering $\Sigma\Pi\Sigma$ formulas, we could study how to convert these formulas to be homogeneous, where each node computes a homogeneous polynomial. To do this, one idea is to apply (1) to each product gate and then compute each resulting elementary symmetric polynomial using homogeneous circuits (for instance, we could use the upper bound of [28] for homogeneous computation of $e_{d}^{n}$ by $\Sigma\Pi\Sigma\Pi$ formulas of size $n2^{O(\sqrt{d})}$, but this only works for characteristic zero fields). This idea can be extended to work for deeper constant-depth circuits.
Fields of positive characteristic
Specifically, in this paper, we will study such computations over fields of positive characteristic. Recall that the characteristic of a field is the smallest $k$ such that adding one to itself $k$ times gives zero, where we say that characteristic is zero if no such $k$ exists. For a positive prime $p>1$, the easiest example of a field of characteristic $p$ is the integers modulo $p$, or $\mathbb{F}_{p}=\{0,\ldots,p-1\}$. Fields of characteristic zero are generally seen as more “natural,” as examples include the real and complex numbers. Specifically, if results require division by natural numbers, they often fail in certain characteristics. For this reason, many results in algebraic complexity theory only work in fields of characteristic zero. This paper is focused on extending some of these results to fields of arbitrary characteristic, but it is also natural to ask why we should study this case.
One simple reasoning follows from the underlying algebraic complexity theory program, whose major goal is to show that VP is not equal to VNP (also known as Valiant’s hypothesis). From [5], we see that it is sufficient to show that VP equals VNP in a field of any characteristic to collapse the polynomial hierarchy. In fact, classical complexity problems are usually posed over alphabets described by $\{0,1\}$, which can be viewed as a field of characteristic two. It is thus not only interesting to study positive characteristics, but fundamentally, our strategy to approach this problem should be independent of the underlying field.
Interestingly, the issue of positive characteristic also shows up in the field of proof complexity, whose focus is on developing lower bounds for the size of “proofs” of logical statements in various proof systems. Specifically, one major proof systems studied is known as the $\text{AC}^{0}[p]$-Frege proof systems, which are bounded-depth proof systems with $\operatorname{MOD}_{p}$ gates, where a major open problem is to find superpolynomial lower bounds for this system. The theory of proof complexity was brought into the world of algebraic complexity through the introduction of the Ideal Proof System (IPS), which defines its proofs using algebraic circuits. In the influential work [17], it was shown that superpolynomial lower bounds for IPS over fields of characteristic $p\neq 0$ give superpolynomial $\text{AC}^{0}[p]$-Frege lower bounds. They further showed that superpolynomial circuit lower bounds lead to superpolynomial IPS lower bounds, so finding such lower bounds over fields of nonzero characteristic can lead to solving this open problem in proof complexity.
Another field containing interesting applications comes from polynomial identity testing (PIT). We typically define the problem of PIT by, given a polynomial belonging to some class of polynomials in a predefined representation, can you write a deterministic algorithm to determine if it is identically the zero polynomial. One major line of research in algebraic complexity is finding polynomial-time PIT algorithms for interesting subclasses of polynomials. In the search of such algorithms, the relationship between lower bounds and PIT algorithms has been uncovered, revealing that lower bounds often lead to PIT algorithms. For instance, in [19], it was shown that exponential lower bounds for the permanent would lead to polynomial-time PIT algorithms for polynomial-sized circuits. Unfortunately, if the lower bounds discovered only apply for fields of characteristic zero, then so will the PIT algorithms. If we would like our PIT algorithms to work over any field, we also need our lower bound strategies to work in these cases.
Recently, there has been a theme of extending lower bound results from fields of characteristic zero to fields of positive characteristic. This trend can be seen in papers such as [2], [8], [11], [3], and [4].
Border complexity
For some of our results, we will focus on studying the approximate complexity of computing polynomials, known as border complexity. To define border complexity, we introduce a new variable $\epsilon$ and extend our underlying field $\mathbb{F}$ to the field of fractions $\mathbb{F}(\epsilon)$. Then, instead of requiring our computational model to compute a polynomial $f$ exactly, it must compute some $\epsilon^{N}f(x)+\epsilon^{N+1}Q(x,\epsilon)$ for some $N\in\mathbb{N}$ and $Q\in\mathbb{F}[x,\epsilon]$. The idea of border complexity is well-studied through the geometric complexity theory program, which hopes to use techniques from algebraic geometry and representation theory to give border complexity lower bounds. Perhaps the most fundamentally interesting reason to study border complexity is that most of our lower bound techniques also work in the world of border complexity, meaning that our techniques do not separate the cases of exact and approximate computation. In spite of this, we know that the border setting can add a lot of power; for example, it was shown in [22] that, in fields of characteristic zero, all polynomials can be approximated by the sum of two (very large) products of affine forms, whereas exact computation can require $\Omega(n)$ lower bounds. Thus, a deeper understanding of the power difference between exact and approximate computation is fundamental to the study of algebraic complexity.
1.1 Results
@@
We have considered computation of the elementary symmetric polynomials using formulas over fields of positive characteristic. By the Ben-Or construction, we know that there is a $\Sigma\Pi\Sigma$ circuit computing $e_{d}^{n}$ of size $O(n^{2})$ over infinite fields (independent of characteristic). In this paper, we extend the results of [6] and show that their results on fields of zero characteristic extend to fields of positive characteristic, namely that this upper bound is tight for formulas for certain elementary symmetric polynomials.
@@
In the second part of the paper, we show that the results of [29] and [22] do not extend to fields of positive characteristic. First, we show that the symmetric model is not universal over such fields. In fact, we show something stronger; we show that there are some polynomials that cannot be represented by $k$-linear combinations of projections of the symmetric model, even if we extend to the border setting.
@@
If $\operatorname{char}(\mathbb{F})\neq 0$, then, for every $n\in\mathbb{N}$, there is a homogeneous polynomial $f\in\mathbb{F}[x_{1},\ldots,x_{n}]$ of degree $d$ such that, for every linear $L_{1}^{(1)},\ldots,L_{m_{1}}^{(1)},L_{1}^{(2)},\ldots,L_{m_{k}}^{(k)}\in\mathbb{F}[x_{1},\ldots,x_{n}]$ and constants $c_{1},\ldots,c_{k}\in\mathbb{F}$,
$f\neq\sum_{i=1}^{k}c_{i}e_{d}^{m_{i}}(L_{1}^{(i)},\ldots,L_{m_{i}}^{(i)}),$ (2)
for $k=o(n)$. Further, this applies in the border setting.
As we have previously mentioned, the symmetric model is intimately related to the models studied in [22], namely the $\overline{\Sigma^{[k]}\Pi\Sigma}$ model. Through this, as an immediate corollary of Theorem 1.2, we show that the results of [22] also do not extend to fields of positive characteristic. In fact, we prove, under such bounds, $\Omega(n)$ lower bounds for the border affine Chow rank of certain polynomials.
@@
If $\operatorname{char}(\mathbb{F})\neq 0$, then there exists a polynomial in at most $n$ variables that is not in $\overline{\Sigma^{[k]}\Pi\Sigma}$ for $k=o(n)$.
1.2 Proof techniques
Proof of Theorem 1.1
We will first focus on proving Theorem 1.1 in Section 2. This proof will largely rely on the proof for the case of characteristic zero from [6] but with a modification to the step that relies on the field characteristic. Namely, the proof revolves around studying what we will call the set of order-two zeros of a polynomial.
Algebraic complexity theory focuses on using the algebraic properties of polynomials to split them into classes based on how “hard” they are to compute. One such property that has been utilized (see [14], [21], [6], [20], [1]) is what we will call the order-$2$ zero space. Given a polynomial $f\in\mathbb{F}[x_{1},\ldots,x_{n}]$, we define its order-$2$ zero space, denoted $V_{2}(f)$, to be the points where it vanishes along with its first-order partial derivatives, i.e.,
@@ 1.2 Proof techniques
This idea of the order-$2$ zero space is utilized in [21] and [6] without a name. Then, in [20], it is introduced as the “singular locus” of a polynomial, denoted $\operatorname{sing}(f)$, which is a well-known object in algebraic geometry, in a nod to the notation of [14]. For this paper, we have decided to change this notation, as this definition of singular locus does not precisely align with the algebraic geometric definition. Although it is true that, when $f$ is square-free, this definition exactly characterizes the singular locus ($V_{2}(f)=\operatorname{sing}(f)$), this is not true in general. This follows from the fact that the singular locus of a polynomial is a property of its corresponding hypersurface, a purely geometric object, while the order-$2$ zero space is a property of the polynomial, itself. Specifically, if we consider the power polynomial $p_{d}^{n}=x_{1}^{d}+\dots+x_{n}^{d}$, it is well-known that $\operatorname{sing}(p_{d}^{n})=\{0\}$ (see Example 10.21 of [15]), but, if we consider $\operatorname{char}(\mathbb{F})=q\neq 0$, we trivially observe that $V_{2}(p_{q}^{n})=V(p_{q}^{n})$. We find that letting this definition differ from its natural geometric definition is confusing, so we have, therefore, decided to change it.
@@
Lemma 1.5 (See Lemma 1.7 of [21] and Lemma 3.4 of [6]).
@@
where the first inequality comes from Lemma 5.8 of [6] and the last inequality is a basic fact from algebraic geometry. ∎
@@
The key observation for the proof in [6] was that small formulas computed polynomials with “many” order-$2$ zeros. Then, they prove an upper bound of the dimension of the set of order-$2$ zeros of the elementary symmetric polynomials and utilize this to show an $\Omega(d(n-d))$ lower bounds on formulas that compute them over fields of zero characteristic (with $n$ being the number of variables and $d$ being the degree). For certain selections of $n$ and $d$, this bound is tight, as the Ben-Or construction shows an $O(n^{2})$ upper bound on the size of $\Sigma\Pi\Sigma$ formulas computing $e_{d}^{n}$. This was known to be tight for $\Sigma\Pi\Sigma$ formulas over fields of characteristic zero from [29].
Tight bound on the order-$2$ space of the elementary symmetric polynomials
It was shown in [26], [23], and [6] that $\dim V_{2}(e_{d}^{n})=d-2$ over fields of characteristic zero, where dimension is, of course, defined in terms of affine varieties. No such equivalent statement was known (to the author’s knowledge) for fields of positive characteristic. In this paper, we show that the dimension is slightly different over such fields; specifically, it can vary between being d2d-2 or d1d-1.
@@
For $\operatorname{char}(\mathbb{F})=p$, we have that $d-2\leq\dim V_{2}(e_{d}^{n})\leq d-1$. If we have that
  • $p\neq 0$,
  • $n-d+1\equiv 0\mod p^{\lceil\log_{p}{d}\rceil}$, and
  • $n\geq 2d-1$,
then $\dim V_{2}(e_{d}^{n})=d-1$. On the other hand, if we have that
  • $p=0$,
  • $n-d+1\equiv 0\mod p$, or
  • n<2d1n<2d-1,
then $\dim V_{2}(e_{d}^{n})=d-2$.
Then, for the rest of the proof in [6] to work, we need only that $\dim V_{2}(e_{d}^{n})\leq d-1$, so this lemma suffices for proving the main result.
Interestingly, Lemma 1.7 allows us to also extend the lower bound techniques in [29] used to show that $e_{d}^{n}$ require $\Sigma\Pi\Sigma$ formulas of size $\Omega(d(n-d))$ to fields of positive characteristics [^2] Our results, of course, immediately imply this lower bound, but it is still of interest that we can also extend the results of their main technical lemma. In [29], they show that, over fields of characteristic zero, for every vector space $V\subseteq\mathbb{F}^{n}$ such that $e_{d}^{n}$ vanishes on $V$, we have that $\dim(V)<\frac{n+d}{2}$. For fields of positive characteristic, observe that if we combine Proposition 6 from [16] with Lemma 1.7, we conclude that we have the upper bound of $\dim(V)\leq\frac{n+d-1}{2}$ [^3] Using the fact that we can represent a polynomial as a “strength” decomposition of size $\operatorname{codim}(V)$.
Proof of Theorem 1.3
@@
Then, we clearly have a product decomposition, so we can use Lemma 1.5 to conclude that $\dim V_{2}(f)\geq n-\ell\cdot 2^{d+1}$. Then, if we select $d$ as a constant, we conclude that $\ell\geq\Omega(n-\dim V_{2}(f))$. We should observe that this argument only works in the non-border case, and it would require a careful argument to show that it works in the border setting.